# Teon: Tensorized Orthonormalization Beyond Layer-Wise Muon for Large Language Model Pre-Training

### v1

@@ 3.1 From Layer-wise to Tensor-wise Modeling: The Teon Optimizer
We now introduce Teon, a generalization of the Muon optimizer that extends layer-wise gradient orthogonalization to a tensor-based formulation, which attempts to capture correlations among gradients from multiple layers. A natural approach is to stack a set of layer gradients together into a single tensor and use tensor orthogonalization to constrain the update direction. Specifically, Teon first stack gradients across multiple identical layers $\mathbf{G}^{(1)},\dots,\mathbf{G}^{(K)},\mathbf{G}^{(k)}\in\mathbb{R}^{m\times n}$ into an order-$3$ tensor:
@@ 4.2 Sketch of Proofs
The assumption of rank-$1$ and shared singular vector in Lemma 4.8 characterizes an extreme case that achieves $\sqrt{K}\times$ improvement. In realistic training, layer-wise gradients are not rank-$1$ but typically low-rank (Gur-Ari et al., 2018; Zhao et al., 2024). In this case, Teon achieves a gain that is smaller than but close to $\sqrt{K}\times$, with the magnitude controlled by the alignment of the leading singular vectors across layers. This motivates the Proposition 4.6: stacking layers whose gradients exhibit strong top-singular-vector alignment as a practical criterion for applying cross-layer (mode-1 / mode-2) orthogonalization.
@@ A.1 Tensor norms for Muon and Teon
Throughout, let $\bm{\mathcal{X}}\in\mathbb{R}^{m\times n\times K}$ with slices $\bm{\mathcal{X}}[:,:,k]=X^{(k)}\in\mathbb{R}^{m\times n}$ for $k\in[K]$, and let $\mathcal{M}_{i}(\cdot)$ denote the mode-$i$ matricization.
@@
$\|\bm{\mathcal{X}}\|_{\textsc{muon}}\;:=\;\max_{k\in[K]}\|X^{(k)}\|_{\mathrm{op}},$ (14)
@@
$\|\bm{\mathcal{X}}\|_{\textsc{muon},*}\;:=\;\sum_{k=1}^{K}\|X^{(k)}\|_{*}.$ (15)
@@
$=\sup_{\|Y\|_{\mathrm{op}}\leq 1}\ \langle\mathcal{M}_{i}(\bm{\mathcal{X}}),Y\rangle_{F}=\|\mathcal{M}_{i}(\bm{\mathcal{X}})\|_{*},$ (17)
@@
$\big(\|\cdot\|_{\textsc{muon}}\big)_{*}(\bm{\mathcal{X}})\;=\;\sum_{k=1}^{K}\|X^{(k)}\|_{*}.$ (18)
@@
$\big(\|\cdot\|_{\textsc{muon}}\big)_{*}(\bm{\mathcal{X}})$ $=\sup_{\|\bm{\mathcal{Y}}\|_{\textsc{muon}}\leq 1}\ \langle\bm{\mathcal{X}},\bm{\mathcal{Y}}\rangle=\sup_{\max_{k}\|Y^{(k)}\|_{\mathrm{op}}\leq 1}\sum_{k=1}^{K}\langle X^{(k)},Y^{(k)}\rangle_{F}$
$=\sum_{k=1}^{K}\sup_{\|Y^{(k)}\|_{\mathrm{op}}\leq 1}\ \langle X^{(k)},Y^{(k)}\rangle_{F}=\sum_{k=1}^{K}\|X^{(k)}\|_{*},$ (19)
@@
Let $\bm{\mathcal{X}}\in\mathbb{R}^{A\times B\times K}$ with slices $\bm{\mathcal{X}}[:,:,k]=X^{(k)}$.
@@
$\|\bm{\mathcal{X}}\|_{\textsc{muon}}\;:=\;\max_{k\in[K]}\|X^{(k)}\|_{\mathrm{op}}.$
For each slice, $\|X^{(k)}\|_{\mathrm{op}}\leq\|X^{(k)}\|_{F}$. Hence,
$\|\bm{\mathcal{X}}\|_{\textsc{muon}}\leq\max_{k}\|X^{(k)}\|_{F}\leq\Big(\sum_{k=1}^{K}\|X^{(k)}\|_{F}^{2}\Big)^{1/2}=\|\bm{\mathcal{X}}\|_{F},$
@@ A.2 Teon/Muon step as steepest descent
where $\|\bm{\mathcal{X}}\|_{\textsc{teon}-i}:=\|\mathcal{M}_{i}(\bm{\mathcal{X}})\|_{\mathrm{op}}$. Let $\mathcal{M}_{i}(\bm{\mathcal{G}})=U\Sigma V^{\top}$ be a (thin) SVD. Then an optimal solution is
$\Delta\bm{\mathcal{W}}\;=\;-\eta\ \mathcal{M}_{i}^{-1}\!\big(UV^{\top}\big).$ (21)
@@
$=\mathcal{M}_{i}^{-1}\!\left(\arg\min_{\|Y\|_{\mathrm{op}}\leq\eta}\left\langle\mathcal{M}_{i}(\bm{\mathcal{G}}),\ Y\right\rangle_{F}\right).$ (23)
Let $Y=\eta\tilde{Y}$ with $\|\tilde{Y}\|_{\mathrm{op}}\leq 1$. Then
$\arg\min_{\|Y\|_{\mathrm{op}}\leq\eta}\left\langle\mathcal{M}_{i}(\bm{\mathcal{G}}),\ Y\right\rangle_{F}$ $=-\eta\ \arg\max_{\|\tilde{Y}\|_{\mathrm{op}}\leq 1}\left\langle\mathcal{M}_{i}(\bm{\mathcal{G}}),\ \tilde{Y}\right\rangle_{F}.$ (24)
If $\mathcal{M}_{i}(\bm{\mathcal{G}})=U\Sigma V^{\top}$ is a thin SVD, then the maximizer of $\max_{\|\tilde{Y}\|_{\mathrm{op}}\leq 1}\langle\mathcal{M}_{i}(\bm{\mathcal{G}}),\tilde{Y}\rangle_{F}$ is $\tilde{Y}^{\star}=UV^{\top}$ (a standard duality fact: the dual of $\|\cdot\|_{\mathrm{op}}$ is $\|\cdot\|_{*}$). Substituting $\tilde{Y}^{\star}$ into (24) and then into (23)–(22) yields $\Delta\bm{\mathcal{W}}=-\eta\,\mathcal{M}_{i}^{-1}(UV^{\top})$. ∎
@@
where the Muon norm is $\|\bm{\mathcal{X}}\|_{\textsc{muon}}:=\max_{k\in[K]}\|\bm{\mathcal{X}}[:,:,k]\|_{\mathrm{op}}$. For each slice, let $\bm{\mathcal{G}}[:,:,k]=G^{(k)}=U^{(k)}\Sigma^{(k)}(V^{(k)})^{\top}$ be a (thin) SVD. Then an optimal solution has slices
$\Delta\bm{\mathcal{W}}[:,:,k]\;=\;-\eta\ U^{(k)}(V^{(k)})^{\top},\qquad\forall k\in[K].$ (26)
@@
By definition of $\|\cdot\|_{\textsc{muon}}$, the constraint $\|\Delta\bm{\mathcal{W}}\|_{\textsc{muon}}\leq\eta$ is equivalent to $\|\Delta\bm{\mathcal{W}}[:,:,k]\|_{\mathrm{op}}\leq\eta$ for all $k\in[K]$. Moreover, $\langle\bm{\mathcal{G}},\Delta\bm{\mathcal{W}}\rangle=\sum_{k=1}^{K}\langle G^{(k)},\Delta W^{(k)}\rangle_{F}$. Therefore (25) decouples across layers:
$\min_{\|\Delta\bm{\mathcal{W}}\|_{\textsc{muon}}\leq\eta}\ \langle\bm{\mathcal{G}},\Delta\bm{\mathcal{W}}\rangle$ $=\sum_{k=1}^{K}\ \min_{\|\Delta W^{(k)}\|_{\mathrm{op}}\leq\eta}\ \langle G^{(k)},\Delta W^{(k)}\rangle_{F}.$ (27)
Each summand is the matrix trust-region subproblem under $\|\cdot\|_{\mathrm{op}}$. Let $G^{(k)}=U^{(k)}\Sigma^{(k)}(V^{(k)})^{\top}$ be a thin SVD. By the same duality argument as in Proposition A.6, the minimizer is $\Delta W^{(k)}=-\eta\,U^{(k)}(V^{(k)})^{\top}$. Stacking these slices yields (26). ∎
@@ A.3 Proof for comparability of the norm and dual norm of teon and muon
Write $\mathcal{M}_{1}(\bm{\mathcal{X}})=[X^{(1)},\dots,X^{(K)}]\in\mathbb{R}^{A\times(BK)}$.
Lower bound. Fix $k\in[K]$. For any unit vector $v\in\mathbb{R}^{B}$, let $z\in\mathbb{R}^{BK}$ be the unit block vector whose $k$-th block equals $v$ and all other blocks are zero. Then $\mathcal{M}_{1}(\bm{\mathcal{X}})z=X^{(k)}v$, hence
$\|\mathcal{M}_{1}(\bm{\mathcal{X}})\|_{\mathrm{op}}=\sup_{\|z\|_{2}=1}\|\mathcal{M}_{1}(\bm{\mathcal{X}})z\|_{2}\geq\sup_{\|v\|_{2}=1}\|X^{(k)}v\|_{2}=\|X^{(k)}\|_{\mathrm{op}}.$
@@
Upper bound. Let $z\in\mathbb{R}^{BK}$ be any unit vector with blocks $z=[z_{1}^{\top}~\cdots~z_{K}^{\top}]^{\top}$, $z_{k}\in\mathbb{R}^{B}$. Then $\mathcal{M}_{1}(\bm{\mathcal{X}})z=\sum_{k=1}^{K}X^{(k)}z_{k}$, and
$\|\mathcal{M}_{1}(\bm{\mathcal{X}})z\|_{2}$ $\leq\sum_{k=1}^{K}\|X^{(k)}z_{k}\|_{2}\leq\sum_{k=1}^{K}\|X^{(k)}\|_{\mathrm{op}}\,\|z_{k}\|_{2}$
$\leq\Big(\sum_{k=1}^{K}\|X^{(k)}\|_{\mathrm{op}}^{2}\Big)^{1/2}\Big(\sum_{k=1}^{K}\|z_{k}\|_{2}^{2}\Big)^{1/2}=\Big(\sum_{k=1}^{K}\|X^{(k)}\|_{\mathrm{op}}^{2}\Big)^{1/2}.$
Taking $\sup_{\|z\|_{2}=1}$ yields
$\|\mathcal{M}_{1}(\bm{\mathcal{X}})\|_{\mathrm{op}}\leq\Big(\sum_{k=1}^{K}\|X^{(k)}\|_{\mathrm{op}}^{2}\Big)^{1/2}\leq\sqrt{K}\,\max_{k}\|X^{(k)}\|_{\mathrm{op}}=\sqrt{K}\ \|\bm{\mathcal{X}}\|_{\textsc{muon}}.$
@@
Under the standard mode-2 unfolding, $\mathcal{M}_{2}(\bm{\mathcal{X}})=[(X^{(1)})^{\top}~\cdots~(X^{(K)})^{\top}]\in\mathbb{R}^{B\times(AK)}$, which is again a horizontal concatenation of $K$ blocks. The proof is identical to Lemma A.8, using $\|(X^{(k)})^{\top}\|_{\mathrm{op}}=\|X^{(k)}\|_{\mathrm{op}}$. ∎
@@
Let $\|\cdot\|_{a}$ and $\|\cdot\|_{b}$ be norms on an inner-product space with duals $\|\cdot\|_{a,*}$ and $\|\cdot\|_{b,*}$. If there exist $\alpha,\beta>0$ such that for all $x$,
$\alpha\|x\|_{a}\leq\|x\|_{b}\leq\beta\|x\|_{a},$
then for all $g$,
$\frac{1}{\beta}\|g\|_{a,*}\leq\|g\|_{b,*}\leq\frac{1}{\alpha}\|g\|_{a,*}.$

### v2

@@ 3.1 From Layer-wise to Tensor-wise Modeling: The Teon Optimizer
We now introduce Teon, a generalization of the Muon optimizer that extends layer-wise gradient orthogonalization to a tensor-based formulation, which attempts to capture correlations among gradients from multiple layers. A natural approach is to stack a set of layer gradients together into a single tensor and use tensor orthogonalization to constrain the update direction. Specifically, Teon first stack gradients across multiple identical layers $\mathbf{G}^{(1)},\dots,\mathbf{G}^{(K)},\mathbf{G}^{(k)}\in\mathbb{R}^{m\times n}$ into an order-33 tensor:
@@ 4.2 Sketch of Proofs
The assumption of rank-11 and shared singular vector in Lemma 4.8 characterizes an extreme case that achieves $\sqrt{K}\times$ improvement. In realistic training, layer-wise gradients are not rank-11 but typically low-rank (Gur-Ari et al., 2018; Zhao et al., 2024). In this case, Teon achieves a gain that is smaller than but close to $\sqrt{K}\times$, with the magnitude controlled by the alignment of the leading singular vectors across layers. This motivates the Proposition 4.6: stacking layers whose gradients exhibit strong top-singular-vector alignment as a practical criterion for applying cross-layer (mode-1 / mode-2) orthogonalization.
@@ A.1 Tensor norms for Muon and Teon
Throughout, let $\bm{\mathcal{X}}\in\mathbb{R}^{m\times n\times K}$ with slices $\bm{\mathcal{X}}[:,:,k]=\mathbf{X}^{(k)}\in\mathbb{R}^{m\times n}$ for $k\in[K]$, and let $\mathcal{M}_{i}(\cdot)$ denote the mode-$i$ matricization.
@@
$\|\bm{\mathcal{X}}\|_{\textsc{muon}}\;:=\;\max_{k\in[K]}\|\mathbf{X}^{(k)}\|_{\mathrm{op}},$ (14)
@@
$\|\bm{\mathcal{X}}\|_{\textsc{muon},*}\;:=\;\sum_{k=1}^{K}\|\mathbf{X}^{(k)}\|_{*}.$ (15)
@@
$=\sup_{\|\mathbf{Y}\|_{\mathrm{op}}\leq 1}\ \langle\mathcal{M}_{i}(\bm{\mathcal{X}}),\mathbf{Y}\rangle_{F}=\|\mathcal{M}_{i}(\bm{\mathcal{X}})\|_{*},$ (17)
@@
$\big(\|\cdot\|_{\textsc{muon}}\big)_{*}(\bm{\mathcal{X}})\;=\;\sum_{k=1}^{K}\|\mathbf{X}^{(k)}\|_{*}.$ (18)
@@
$\big(\|\cdot\|_{\textsc{muon}}\big)_{*}(\bm{\mathcal{X}})$ $=\sup_{\|\bm{\mathcal{Y}}\|_{\textsc{muon}}\leq 1}\ \langle\bm{\mathcal{X}},\bm{\mathcal{Y}}\rangle=\sup_{\max_{k}\|\mathbf{Y}^{(k)}\|_{\mathrm{op}}\leq 1}\sum_{k=1}^{K}\langle\mathbf{X}^{(k)},\mathbf{Y}^{(k)}\rangle_{F}$
$=\sum_{k=1}^{K}\sup_{\|\mathbf{Y}^{(k)}\|_{\mathrm{op}}\leq 1}\ \langle\mathbf{X}^{(k)},\mathbf{Y}^{(k)}\rangle_{F}=\sum_{k=1}^{K}\|\mathbf{X}^{(k)}\|_{*},$ (19)
@@
Let $\bm{\mathcal{X}}\in\mathbb{R}^{A\times B\times K}$ with slices $\bm{\mathcal{X}}[:,:,k]=\mathbf{X}^{(k)}$.
@@
$\|\bm{\mathcal{X}}\|_{\textsc{muon}}\;:=\;\max_{k\in[K]}\|\mathbf{X}^{(k)}\|_{\mathrm{op}}.$
For each slice, $\|\mathbf{X}^{(k)}\|_{\mathrm{op}}\leq\|\mathbf{X}^{(k)}\|_{F}$. Hence,
$\|\bm{\mathcal{X}}\|_{\textsc{muon}}\leq\max_{k}\|\mathbf{X}^{(k)}\|_{F}\leq\Big(\sum_{k=1}^{K}\|\mathbf{X}^{(k)}\|_{F}^{2}\Big)^{1/2}=\|\bm{\mathcal{X}}\|_{F},$
@@ A.2 Teon/Muon step as steepest descent
where $\|\bm{\mathcal{X}}\|_{\textsc{teon}-i}:=\|\mathcal{M}_{i}(\bm{\mathcal{X}})\|_{\mathrm{op}}$. Let $\mathcal{M}_{i}(\bm{\mathcal{G}})=\mathbf{U}\mathbf{\Sigma}\mathbf{V}^{\top}$ be a (thin) SVD. Then an optimal solution is
$\Delta\bm{\mathcal{W}}\;=\;-\eta\ \mathcal{M}_{i}^{-1}\!\big(\mathbf{U}\mathbf{V}^{\top}\big).$ (21)
@@
$=\mathcal{M}_{i}^{-1}\!\left(\arg\min_{\|\mathbf{Y}\|_{\mathrm{op}}\leq\eta}\left\langle\mathcal{M}_{i}(\bm{\mathcal{G}}),\ \mathbf{Y}\right\rangle_{F}\right).$ (23)
Let $\mathbf{Y}=\eta\tilde{\mathbf{Y}}$ with $\|\tilde{\mathbf{Y}}\|_{\mathrm{op}}\leq 1$. Then
$\arg\min_{\|\mathbf{Y}\|_{\mathrm{op}}\leq\eta}\left\langle\mathcal{M}_{i}(\bm{\mathcal{G}}),\ \mathbf{Y}\right\rangle_{F}$ $=-\eta\ \arg\max_{\|\tilde{\mathbf{Y}}\|_{\mathrm{op}}\leq 1}\left\langle\mathcal{M}_{i}(\bm{\mathcal{G}}),\ \tilde{\mathbf{Y}}\right\rangle_{F}.$ (24)
If $\mathcal{M}_{i}(\bm{\mathcal{G}})=\mathbf{U}\mathbf{\Sigma}\mathbf{V}^{\top}$ is a thin SVD, then the maximizer of $\max_{\|\tilde{\mathbf{Y}}\|_{\mathrm{op}}\leq 1}\langle\mathcal{M}_{i}(\bm{\mathcal{G}}),\tilde{\mathbf{Y}}\rangle_{F}$ is $\tilde{\mathbf{Y}}^{\star}=\mathbf{U}\mathbf{V}^{\top}$ (a standard duality fact: the dual of $\|\cdot\|_{\mathrm{op}}$ is $\|\cdot\|_{*}$). Substituting $\tilde{\mathbf{Y}}^{\star}$ into (24) and then into (23)–(22) yields $\Delta\bm{\mathcal{W}}=-\eta\,\mathcal{M}_{i}^{-1}(\mathbf{U}\mathbf{V}^{\top})$. ∎
@@
where the Muon norm is $\|\bm{\mathcal{X}}\|_{\textsc{muon}}:=\max_{k\in[K]}\|\bm{\mathcal{X}}[:,:,k]\|_{\mathrm{op}}$. For each slice, let $\bm{\mathcal{G}}[:,:,k]=G^{(k)}=\mathbf{U}^{(k)}\mathbf{\Sigma}^{(k)}(\mathbf{V}^{(k)})^{\top}$ be a (thin) SVD. Then an optimal solution has slices
$\Delta\bm{\mathcal{W}}[:,:,k]\;=\;-\eta\ \mathbf{U}^{(k)}(\mathbf{V}^{(k)})^{\top},\qquad\forall k\in[K].$ (26)
@@
By definition of $\|\cdot\|_{\textsc{muon}}$, the constraint $\|\Delta\bm{\mathcal{W}}\|_{\textsc{muon}}\leq\eta$ is equivalent to $\|\Delta\bm{\mathcal{W}}[:,:,k]\|_{\mathrm{op}}\leq\eta$ for all $k\in[K]$. Moreover, $\langle\bm{\mathcal{G}},\Delta\bm{\mathcal{W}}\rangle=\sum_{k=1}^{K}\langle\mathbf{G}^{(k)},\Delta\mathbf{W}^{(k)}\rangle_{F}$. Therefore (25) decouples across layers:
$\min_{\|\Delta\bm{\mathcal{W}}\|_{\textsc{muon}}\leq\eta}\ \langle\bm{\mathcal{G}},\Delta\bm{\mathcal{W}}\rangle$ $=\sum_{k=1}^{K}\ \min_{\|\Delta\mathbf{W}^{(k)}\|_{\mathrm{op}}\leq\eta}\ \langle\mathbf{G}^{(k)},\Delta\mathbf{W}^{(k)}\rangle_{F}.$ (27)
Each summand is the matrix trust-region subproblem under $\|\cdot\|_{\mathrm{op}}$. Let $G^{(k)}=\mathbf{U}^{(k)}\mathbf{\Sigma}^{(k)}(\mathbf{V}^{(k)})^{\top}$ be a thin SVD. By the same duality argument as in Proposition A.6, the minimizer is $\Delta\mathbf{W}^{(k)}=-\eta\,\mathbf{U}^{(k)}(\mathbf{V}^{(k)})^{\top}$. Stacking these slices yields (26). ∎
@@ A.3 Proof for comparability of the norm and dual norm of teon and muon
Write $\mathcal{M}_{1}(\bm{\mathcal{X}})=[\mathbf{X}^{(1)},\dots,\mathbf{X}^{(K)}]\in\mathbb{R}^{A\times(BK)}$.
Lower bound. Fix $k\in[K]$. For any unit vector $\mathbf{v}\in\mathbb{R}^{B}$, let $\mathbf{z}\in\mathbb{R}^{BK}$ be the unit block vector whose $k$-th block equals $\mathbf{v}$ and all other blocks are zero. Then $\mathcal{M}_{1}(\bm{\mathcal{X}})\mathbf{z}=\mathbf{X}^{(k)}\mathbf{v}$, hence
$\|\mathcal{M}_{1}(\bm{\mathcal{X}})\|_{\mathrm{op}}=\sup_{\|\mathbf{z}\|_{2}=1}\|\mathcal{M}_{1}(\bm{\mathcal{X}})\mathbf{z}\|_{2}\geq\sup_{\|\mathbf{v}\|_{2}=1}\|\mathbf{X}^{(k)}\mathbf{v}\|_{2}=\|\mathbf{X}^{(k)}\|_{\mathrm{op}}.$
@@
Upper bound. Let $\mathbf{z}\in\mathbb{R}^{BK}$ be any unit vector with blocks $\mathbf{z}=[\mathbf{z}_{1}^{\top}~\cdots~\mathbf{z}_{K}^{\top}]^{\top}$, $\mathbf{z}_{k}\in\mathbb{R}^{B}$. Then $\mathcal{M}_{1}(\bm{\mathcal{X}})\mathbf{z}=\sum_{k=1}^{K}\mathbf{X}^{(k)}\mathbf{z}_{k}$, and
$\|\mathcal{M}_{1}(\bm{\mathcal{X}})\mathbf{z}\|_{2}$ $\leq\sum_{k=1}^{K}\|\mathbf{X}^{(k)}\mathbf{z}_{k}\|_{2}\leq\sum_{k=1}^{K}\|\mathbf{X}^{(k)}\|_{\mathrm{op}}\,\|\mathbf{z}_{k}\|_{2}$
$\leq\Big(\sum_{k=1}^{K}\|\mathbf{X}^{(k)}\|_{\mathrm{op}}^{2}\Big)^{1/2}\Big(\sum_{k=1}^{K}\|\mathbf{z}_{k}\|_{2}^{2}\Big)^{1/2}=\Big(\sum_{k=1}^{K}\|\mathbf{X}^{(k)}\|_{\mathrm{op}}^{2}\Big)^{1/2}.$
Taking $\sup_{\|\mathbf{z}\|_{2}=1}$ yields
$\|\mathcal{M}_{1}(\bm{\mathcal{X}})\|_{\mathrm{op}}\leq\Big(\sum_{k=1}^{K}\|\mathbf{X}^{(k)}\|_{\mathrm{op}}^{2}\Big)^{1/2}\leq\sqrt{K}\,\max_{k}\|\mathbf{X}^{(k)}\|_{\mathrm{op}}=\sqrt{K}\ \|\bm{\mathcal{X}}\|_{\textsc{muon}}.$
@@
Under the standard mode-2 unfolding, $\mathcal{M}_{2}(\bm{\mathcal{X}})=[(\mathbf{X}^{(1)})^{\top}~\cdots~(\mathbf{X}^{(K)})^{\top}]\in\mathbb{R}^{B\times(AK)}$, which is again a horizontal concatenation of $K$ blocks. The proof is identical to Lemma A.8, using $\|(\mathbf{X}^{(k)})^{\top}\|_{\mathrm{op}}=\|\mathbf{X}^{(k)}\|_{\mathrm{op}}$. ∎
@@
Let $\|\cdot\|_{a}$ and $\|\cdot\|_{b}$ be norms on an inner-product space with duals $\|\cdot\|_{a,*}$ and $\|\cdot\|_{b,*}$. If there exist $\alpha,\beta>0$ such that for all $\mathbf{x}$,
$\alpha\|\mathbf{x}\|_{a}\leq\|\mathbf{x}\|_{b}\leq\beta\|\mathbf{x}\|_{a},$
then for all $\mathbf{g}$,
$\frac{1}{\beta}\|\mathbf{g}\|_{a,*}\leq\|\mathbf{g}\|_{b,*}\leq\frac{1}{\alpha}\|\mathbf{g}\|_{a,*}.$